An extremely aerodynamic penny is thrown vertically downward off an observation tower. 4 seconds later, it impacts the ground 45m below. Find the impact velocity if g= 9.81m/s^2.
1
Expert's answer
2019-10-17T11:08:05-0400

We have:


"h=v_0t+0.5gt^2"

"45=v_0(4)+0.5(9.81)(4)^2"

"v_0=-8.37\frac{m}{s}"

The direction was upwards.


"v=v_0+gt=-8.37+9.81(4)=30.9\frac{m}{s}"